Please look at two pics, if the answers are right you will get thumbs up, wrong thumbs down. So don’t answer this question till you 100 percent sure. Show me how to do it with the formula.
s. You own a $1,00-par zero-coupon bond that has 5 years of remaining maturity. You plan on selling the bond in one year and believe that the required yield next year will have the following probability distribution: Required Yield B Bond Price Probability 0.3 0.7 7% 10% Note that the required yield can be interpreted as the discount rate in the formula discussed in class. What is your expected required yield when you sell the bond? a. b. Calculate the variance of the required yield. c. Calculate the bonds price in each situation and complete the right column. (Hint: one year later, the bond has 4 years of remaining maturity) What is your expected price when you sell the bond? (This exercise shows an example that the expected price may not be equal to the price of bond with the expected required yield.) d.
4. Discount Bond For any one year discount bond F Face value of the discount bond P current price of the discount bond The yield to maturity equals the increase in price over the year divided by the initial price. As with a coupon bond, the yield to maturity is negatively related to the current bond price.
Answer #1

a. Expected required yield = 0.3*7% + 0.7*10% = 9.10%

b. The variance on the required yield is calculated as shown below:

Probability P Return R R - E R (R - E R)^2 (R - E R)^2*P
0.3 7.00% -0.021 0.000441 0.0001323
0.7 10.00% 0.009 8.1E-05 5.67E-05
Expected Return (E R) 9.10%
Variance 0.000189

Variance = 0.000189

c. The price of the bond in each situation :

At 7%, Price of the bond = FV/(1+r)^n = 1000/(1+0.07)^4 = $762.90

At 10%, Price of the bond = 1000/(1+0.10)^4 = $683.01

Probability Required Yield Bond Price
0.3 7.00% $     762.90
0.7 10.00% $     683.01

d. Expected price is calculating the expected yield in part (a)

Expected price = 1000/(1+0.091)^4 = $705.83